# Measuring the Speed of Gravity and the Cosmic Expansion with Time Delays between Gravity and Light from Binary Neutron Stars

### v1

@@ III.2 Catalogs description
The distributions of detected BNS mergers feature two peaks at $\cos\iota\sim-1$ and $\cos\iota\sim 1$. This is the result of the dependence of the GW SNR and the sGRB intrinsic luminosity on the inclination angle $\iota$ [see Eqs. (20) and (23)], which causes the detection probability to be highest when the BNS merger orbital angular momentum is aligned or anti-aligned with respect to our line of sight (face-on/face-away). While GW detections span all $\cos\iota$ values, no sGRBs are detected within the range $\cos\iota\in(-0.70,0.70)$. This is a consequence of the exponential dependency of the sGRB luminosity on the inclination angle [see Eq. (23)].
We detect both GWs and sGRBs up to luminosity distances $D_{L}\sim 800$ Mpc for O5 and $D_{L}\sim 11000\;\text{Mpc}$ for ET. For the O5 case, sGRBs will be detectable for almost all GW signals nearly face-on ($|\cos\iota|>0.70$). However, for ET, the GW detection reach will greatly surpass the sGRB detection reach and we will be able to detect sGRBs only for face-on binaries below $D_{L}\sim 3000\;\text{Mpc}$. Consequently, about $\sim 45\%$ of BNS detections in O5 are followed by an sGRB observation, and this rate drops to just $\sim 5\%$ for ET (even though in terms of absolute numbers GW-sGRB joint detections in ET will be more abundant) [53]. An increase in luminosity distance detection range correlates with a shift and a broadening in the observed time delay $\Delta t_{d}$ distribution as may be clearly seen by comparing Figs. 4 and 5. This is a consequence of their shared dependency on the merger redshift, $z$ [refer to Eqs. (1) and (2))]. The $D_{L}$ $\Delta t_{d}$ distribution could then be used to extract implicit redshift information from a set of detected time delays and luminosity distances. However, for BNS merger events detected in O5, the observed time delay distribution does not evolve significantly with luminosity distance. In such scenario, the redshift effect is very small and, as we will show later, it will be difficult to measure it.
@@ IV Results
Using the catalogs of GW and sGRB detections, we forecast the precision with which the prompt time delay distribution, speed of gravity and $H_{0}$ can be used thanks to time-delay cosmography. To accomplish this, we use IcaroGW a Python tool developed to infer the population properties of CBCs observed through GWs [39]. More precisely, we employ a Markov Chain Monte Carlo algorithm  [^5] The emcee sampling algorithm [54] from Bilby [55, 56], to sample the population parameters. We consider three inference scenarios:
@@
The priors that we set for the population parameters are summarized in Tab. 1. For the prompt time delay distribution, we use a uniform prior between $[0,5]$ s for the mean $\mu_{\Delta t}$ and a flat-in-log prior between $[10^{-3},10]$ s for the standard deviation $\sigma_{\Delta t}$. The choice to use a flat-in-log prior within this range stems from the different $\sigma_{\Delta t}$ values used for the $\Delta t_{s}$ distribution models: the flat-in-log prior guarantees a uniform distribution across all these orders of magnitude. For the Hubble constant, we choose a uniform prior distribution between $[30,140]\,{\rm km\,s^{-1}\,Mpc^{-1}}$. Finally, for $|\hat{\alpha}_{0}|$ we use a uniform prior with a maximum value dependent on the detection and simulation scenario, spanning between $\sim{10^{-19}}$ and $\sim{10^{-15}}$. The different prior range is chosen so that the posterior on $\hat{\alpha}_{0}$ is fully contained in the prior range. The values of the remaining population parameters are fixed in all scenarios. In the remainder of this work, any references to “uncertainty” and “precision” correspond to the median with symmetric 68.3% credible interval.
@@ IV.1 Scenario I: Speed of gravity and prompt time delays
As an example, in Fig. 6, we present the joint posterior distribution for 100 events detected by ET. The interplay between source-frame time delay parameters and the speed of gravity is evident in the strong correlation between $\hat{\alpha}_{0}$ and $\mu_{\Delta t}$ exhibited by the posterior distribution. Considering the expression of the detected time delay in Eq. (2), a lower $\hat{\alpha}_{0}$ must be compensated by a higher source-frame time delay to maintain the observed time delay value fixed at its detected value. Figure 6 clearly shows that the GW/sGRB prompt time delay distribution and speed of gravity can be measured jointly. In Fig. 7 we summarize the precisions achievable for the population parameters $(\hat{\alpha}_{0},\mu_{\Delta t},\sigma_{\Delta t})$.
@@
Figure 7: Precision at the 68.3% credible interval for Scenario I on the population parameters when we combine more and more joint GW-sGRB detections. Each row of subplots represents a deduced population parameter $(\hat{\alpha}_{0},\mu_{\Delta t},\sigma_{\Delta t})$ while each column represents a distinct configuration of the detector network (O5, ET). Each subplot shows 3 curves; one for each prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]$ s).
@@
In our simulations, we successfully constrained the speed of gravity across all observation counts and prompt time delay models with an uncertainty $\Delta\hat{\alpha}_{0}$ ranging from $\sim 2\cdot 10^{-16}$ down to $\sim 10^{-19}$. These constraints are up to $10^{4}$ times better than those derived from the GW170817 measurement [4]. We also notice that the precision on $\hat{\alpha}_{0}$ depends on the GW detection range and prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]\,$s) as well as the number of detected GW events $N_{\text{obs}}$. We find that a “sharper” prompt time delay $\Delta t_{s}$ distribution, i.e., $\sigma_{\Delta t}$ is 10 (100) times smaller than $\sigma_{\Delta t}=1.7\,$s, decreases the uncertainty of an $\hat{\alpha}_{0}$ measurement by a factor $\sim$ 0.1 (0.035). Meanwhile, we find that using more sensitive detectors will improve the constraints on $\hat{\alpha}_{0}$. Specifically, ET will be able to constrain $\hat{\alpha}_{0}$ 3 times better than O5-like sensitivities. The measurements of the prompt time delay mean $\mu_{\Delta t}$ and standard deviation $\sigma_{\Delta t}$ are constrained with an uncertainty of $\sim 0.6-0.01\,$s and $\sim 0.5-0.005\,$s respectively. Similar to $\hat{\alpha}_{0}$, the precision of these constraints depends on both $N_{\text{obs}}$ and the prompt time delay model scenario. However, $\mu_{\Delta t}$ could only be constrained with a minimum of 100, 10 and 4 observations for a prompt time delay distribution with $\sigma=1.7\,s$, 0.17 s, and 0.017 s, respectively. To constrain $\sigma_{\Delta t}$ at least 10 observations were necessary for $\sigma_{\Delta t}=1.7\,$s, while 4 observations sufficed for both 0.17 s and 0.017 s. The uncertainty of the ($\mu_{\Delta t}$, $\sigma_{\Delta t}$) measurements varies with the “sharpness” of the prompt time delay distribution in the same qualitative way as in the $\hat{\alpha}_{0}$ case. From Fig. 7, we conclude that the constraints on the prompt time delay population parameters are less dependent on the GW detector sensitivities than the $\hat{\alpha}_{0}$ constraints.
@@ IV.2 Scenario II: Cosmology and prompt time delays
Figure 8: Precision at the 68.3% credible interval for Scenario II on the population parameters when we combine more and more joint GW-sGRB detections. Each row of subplots represents a deduced population parameter $(H_{0},\mu_{\Delta t},\sigma_{\Delta t})$ while each column represents a distinct configuration of the detector network (O5, ET). Each subplot shows 3 curves; one for each prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]$ s).
In general, we observe that the constraining power on $H_{0}$ increases as the detector network sensitivity improves and the prompt time delay distribution becomes “sharper” (lower $\sigma_{\Delta t}/\mu_{\Delta t}$). For a GW detector network with O5-like sensitivities, $H_{0}$ can be constrained with a precision $<30\%$ when more than 10 sources will be detected and the prompt time delays have a narrow distribution. If the distribution of prompt time delays is wider, with an O5-like detector network, we will need about a thousand of these sources to reach a precision of $\sim 60\%$ on $H_{0}$. Instead, for ET, which will observe GW events at higher redshifts, we find that $H_{0}$ could be measured already at $\sim 30\%$ precision with 100 sources and a distribution of prompt time delays that is mildly spread ($\sigma_{\Delta t}=0.17\,$s). In the ET case, we obtain that with 100 of these sources, it will be possible to constrain $H_{0}$ at $<10\%$ precision, which would be enough to provide useful hints on the $H_{0}$ tension. In the best case scenario, for $N_{\text{obs}}=100$, ET and $\sigma_{\Delta t}=0.017$ s, we obtain a value for the Hubble constant of $H_{0}=63.69^{+2.81}_{-2.71}{\rm km\,s^{-1}\,Mpc^{-1}}$. For the prompt time delay parameters, by comparing Figs. 8 and Fig. 7, we find that the qualitative behaviour for these two parameters is similar to what was found in Scenario I.
@@ IV.3 Scenario III: Fully agnostic
In this case, we find that it will be difficult to constrain the $H_{0}$ inside the prior range used for the analysis even when considering narrow prompt time delay distributions and a thousand sources. Figure 10 displays the uncertainties on the speed of gravity and prompt time delay parameters that we can constrain in this case.
Figure 10: Precision at the 68.3% credible interval for Scenario III on the population parameters when we combine more and more joint GW-sGRB detections. Each row of subplots represents a deduced population parameter $(\hat{\alpha}_{0},\mu_{\Delta t},\sigma_{\Delta t})$ while each column represents a distinct configuration of the detector network (O5, ET). Each subplot shows 3 curves; one for each prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]$ s). We don’t display $H_{0}$ as we are unable to constrain it for such inference Scenario.
The parameters $\hat{\alpha}_{0}$, $\mu_{\Delta t}$ and $\sigma_{\Delta t}$ are constrained with uncertainty of $\sim 2\cdot 10^{-16}-5\cdot 10^{-19}$, $\sim 0.5-0.005\,$s and $\sim 0.5-0.005\,$s, respectively. In comparison to Scenario I (c.f. Fig. 7), these parameters are less constrained. This is a consequence of the fact that in this case we are also fitting for $H_{0}$, which introduces additional degeneracies in the estimation of prompt time delay distributions.
For $H_{0}$, we find that, for each detector and prompt time delay model, we will not be able to measure the Hubble constant even with thousands of GW detections. This is a consequence of the fact that the determination of $\hat{\alpha}_{0}$ degenerate with the prompt time delay distribution, which is a crucial ingredient to get the implicit redshift information required for $H_{0}$. In other words, $H_{0}$ and the GW propagation speed are correlated with each other. The two parameters are degenerate since different combinations of their values can result in the same observed time delay. We display the correlations among the population parameters in Fig. 11, where we show the joint posterior distribution for 100 events detected by ET.

### v2

@@ III.2 Catalogs description
The distributions of detected BNS mergers feature two peaks at $\cos\iota\sim-1$ and $\cos\iota\sim 1$. This is the result of the dependence of the GW SNR and the sGRB intrinsic luminosity on the inclination angle $\iota$ [see Eqs. (20) and (23)], which causes the detection probability to be highest when the BNS merger orbital angular momentum is aligned or antialigned with respect to our line of sight (face-on/face-away). While GW detections span all $\cos\iota$ values, no sGRBs are detected within the range $\cos\iota\in(-0.70,0.70)$. This is a consequence of the exponential dependency of the sGRB luminosity on the inclination angle [see Eq. (23)].
We detect GWs up to luminosity distances $D_{L}\sim 900$ Mpc for O5 and $D_{L}\sim 15000\;\text{Mpc}$ for ET. For the O5 case, sGRBs will be detectable for almost all GW signals nearly face-on ($|\cos\iota|>0.70$). However, for ET, the GW detection reach will greatly surpass the sGRB detection reach and we will be able to detect sGRBs only for face-on binaries below $D_{L}\sim 3500\;\text{Mpc}$. Consequently, about $\sim 44\%$ of BNS detections in O5 are followed by an sGRB observation, and this rate drops to just $\sim 4\%$ for ET (even though in terms of absolute numbers GW-sGRB joint detections in ET will be more abundant) [54]. An increase in luminosity distance detection range correlates with a shift and a broadening in the observed time delay $\Delta t_{d}$ distribution as may be clearly seen by comparing Figs. 4 and 5. This is a consequence of their shared dependency on the merger redshift, $z$ [refer to Eqs. (1) and (2)]. The $D_{L}$ $\Delta t_{d}$ distribution could then be used to extract implicit redshift information from a set of detected time delays and luminosity distances. However, for BNS merger events detected in O5, the observed time delay distribution does not evolve significantly with luminosity distance. In such scenario, the redshift effect is very small and, as we will show later, it will be difficult to measure it.
@@ IV Results
Using the catalogs of GW and sGRB detections, we forecast the precision with which the prompt time delay distribution, speed of gravity and $H_{0}$ can be constrained thanks to time-delay cosmography. To accomplish this, we use icarogw a Python tool developed to infer the population properties of CBCs observed through GWs [40]. More precisely, we employ a Markov Chain Monte Carlo algorithm  [^5] The emcee sampling algorithm [55] from Bilby [56, 57], to sample the population parameters. We consider three inference scenarios:
@@
The priors that we set for the population parameters are summarized in Table 1. For the prompt time delay distribution, we use a uniform prior between $[0,5]$ s for the mean $\mu_{\Delta t}$ and a flat-in-log prior between $[10^{-3},10]$ s for the standard deviation $\sigma_{\Delta t}$. The choice to use a flat-in-log prior within this range stems from the different $\sigma_{\Delta t}$ values used for the $\Delta t_{s}$ distribution models: the flat-in-log prior guarantees a uniform distribution across all these orders of magnitude. For the Hubble constant, we choose a uniform prior distribution between $[30,140]\,{\rm km\,s^{-1}\,Mpc^{-1}}$. Finally, for $|\hat{\alpha}_{0}|$ we use a uniform prior with a maximum value dependent on the detection and simulation scenario, spanning between $\sim{10^{-19}}$ and $\sim{10^{-15}}$. The different prior range is chosen so that the posterior on $\hat{\alpha}_{0}$ is fully contained in the prior range. The values of the remaining population parameters are fixed in all scenarios. In the remainder of this work, any references to “uncertainty” and “precision” correspond to the median with symmetric 68.3% credible interval.
@@ IV.1 Scenario I: Speed of gravity and prompt time delays
As an example, in Fig. 6, we present the joint posterior distribution for 100 events detected by ET. The interplay between source-frame time delay parameters and the speed of gravity is evident in the strong correlation between $\hat{\alpha}_{0}$ and $\mu_{\Delta t}$ exhibited by the posterior distribution. Considering the expression of the detected time delay in Eq. (2), a lower $\hat{\alpha}_{0}$ must be compensated by a higher source-frame time delay to maintain the observed time delay value fixed at its detected value. Figure 6 clearly shows that the GW-sGRB prompt time delay distribution and speed of gravity can be measured jointly. In Fig. 7 we summarize the precisions achievable for the population parameters $(\hat{\alpha}_{0},\mu_{\Delta t},\sigma_{\Delta t})$.
@@
Figure 7: Precision at the 68.3% credible interval for scenario I on the population parameters when we combine more and more joint GW-sGRB detections. Each row of subplots represents a deduced population parameter $(\hat{\alpha}_{0},\mu_{\Delta t},\sigma_{\Delta t})$. Each subplot shows 6 curves; one for each prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]$ s) and configuration of the detector network (O5, ET).
@@
In our simulations, we successfully constrained the speed of gravity across all observation counts and prompt time delay models with an uncertainty $\Delta\hat{\alpha}_{0}$ ranging from $\sim 2\cdot 10^{-16}$ down to $\sim 10^{-19}$. These constraints are up to $10^{4}$ times better than those derived from the GW170817 measurement [4]. We also notice that the precision on $\hat{\alpha}_{0}$ depends on the GW detection range and prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]\,$s) as well as the number of detected GW events $N_{\text{obs}}$. In Fig.7, for a fixed detector sensitivity and prompt time delay distribution model, the precision in inferring the parameters can decrease from 4 to 10 observations. This behavior arises due to statistical fluctuations, as different event realizations lead to variations in the precision of the inferred parameters. The scatter is more significant with fewer observations. In fact, with a sufficiently large number of events, the uncertainty is expected to scale as $1/\sqrt{N_{\rm obs}}$, reducing the impact of these fluctuations. We find that a “sharper” prompt time delay $\Delta t_{s}$ distribution, i.e., $\sigma_{\Delta t}$ is 10 (100) times smaller than $\sigma_{\Delta t}=1.7\,$s, decreases the uncertainty of an $\hat{\alpha}_{0}$ measurement by a factor $\sim$ 0.1 (0.035). Meanwhile, we find that using more sensitive detectors will improve the constraints on $\hat{\alpha}_{0}$. Specifically, ET will be able to constrain $\hat{\alpha}_{0}$ 3 times better than O5-like sensitivities. The measurements of the prompt time delay mean $\mu_{\Delta t}$ and standard deviation $\sigma_{\Delta t}$ are constrained with an uncertainty of $\sim 0.6-0.015\,$s and $\sim 0.5-0.006\,$s respectively. Similar to $\hat{\alpha}_{0}$, the precision of these constraints depends on both $N_{\text{obs}}$ and the prompt time delay model scenario. However, $\mu_{\Delta t}$ could only be constrained with a minimum of 100, 10 and 4 observations for a prompt time delay distribution with $\sigma=1.7$ s, 0.17 s, and 0.017 s, respectively. To constrain $\sigma_{\Delta t}$ at least 10 observations were necessary for $\sigma_{\Delta t}=1.7\,$s, while 4 observations sufficed for both 0.17 s and 0.017 s. The uncertainty of the ($\mu_{\Delta t}$, $\sigma_{\Delta t}$) measurements varies with the “sharpness” of the prompt time delay distribution in the same qualitative way as in the $\hat{\alpha}_{0}$ case. From Fig. 7, we conclude that the constraints on the prompt time delay population parameters are less dependent on the GW detector sensitivities than the $\hat{\alpha}_{0}$ constraints. This is consistent with the expected error budgets from Eq. (2). Assuming Gaussian errors for $\Delta t_{d}$, and a known prompt time delay distribution, we can see that the expected standard deviation on $\hat{\alpha}_{0}$ is inversely proportional to the lookback time $\text{T}_{l}$. Repeating the same procedure for the prompt time delay but fixing $\hat{\alpha}_{0}=0$, provides a precision on $\Delta t_{\rm s}$ inversely proportional to $1+z$. As the lookback time $\text{T}_{l}$ grows faster than $1+z$, for high redshift values (higher detection ranges), the precision on $\hat{\alpha}_{0}$ will improve faster than the ones on prompt time delay distributions.
@@ IV.2 Scenario II: Cosmology and prompt time delays
Figure 8: Precision at the 68.3% credible interval for scenario II on the population parameters when we combine more and more joint GW-sGRB detections. Each row of subplots represents a deduced population parameter $(H_{0},\mu_{\Delta t},\sigma_{\Delta t})$. Each subplot shows 6 curves; one for each prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]$ s) and configuration of the detector network (O5, ET).
In general, we observe that the constraining power on $H_{0}$ increases as the detector network sensitivity improves and the prompt time delay distribution becomes “sharper” (lower $\sigma_{\Delta t}/\mu_{\Delta t}$). For a GW detector network with O5-like sensitivities, $H_{0}$ can be constrained with a precision $<30\%$ when 100 sources will be detected and the prompt time delays have a narrow distribution ($\sigma_{\Delta t}=0.017\,$s). If the distribution of prompt time delays is wider, with an O5-like detector network, we will not be able to constrain $H_{0}$ even with 100 of these sources. Instead, for ET, which will observe GW events at higher redshifts, we find that $H_{0}$ could be measured already at $\sim 30\%$ precision with 100 sources and a distribution of prompt time delays that is mildly spread ($\sigma_{\Delta t}=0.17\,$s). In the ET case, we obtain that with 100 of these sources, it will be possible to constrain $H_{0}$ at $<10\%$ precision, which would be enough to provide useful hints on the $H_{0}$ tension. In the best case scenario, for $N_{\text{obs}}=100$, ET and $\sigma_{\Delta t}=0.017$ s, we obtain a value for the Hubble constant of $H_{0}=63.69^{+2.81}_{-2.71}{\rm km\,s^{-1}\,Mpc^{-1}}$. For the prompt time delay parameters, by comparing Figs. 8 and Fig. 7, we find that the qualitative behavior for these two parameters is similar to what was found in scenario I.
@@ IV.3 Scenario III: Fully agnostic
In this case, we find that it will be difficult to constrain the $H_{0}$ inside the prior range used for the analysis even when considering narrow prompt time delay distributions and a hundred sources. Figure 10 displays the uncertainties on the speed of gravity and prompt time delay parameters that we can constrain in this case.
Figure 10: Precision at the 68.3% credible interval for scenario III on the population parameters when we combine more and more joint GW-sGRB detections. Each row of subplots represents a deduced population parameter $(\hat{\alpha}_{0},\mu_{\Delta t},\sigma_{\Delta t})$. Each subplot shows 6 curves; one for each prompt time delay model ($\sigma_{\Delta t}=[1.7,0.17,0.017]$ s) and configuration of the detector network (O5, ET). We do not display $H_{0}$ as we are unable to constrain it for such inference scenario.
The parameters $\hat{\alpha}_{0}$, $\mu_{\Delta t}$ and $\sigma_{\Delta t}$ are constrained with uncertainty of $\sim 2\cdot 10^{-16}-10^{-18}$, $\sim 0.5-0.02\,$s and $\sim 0.5-0.006\,$s, respectively. In comparison to scenario I (cf. Fig. 7), these parameters are less constrained. This is a consequence of the fact that in this case we are also fitting for $H_{0}$, which introduces additional degeneracies in the estimation of prompt time delay distributions.
For $H_{0}$, we find that, for each detector and prompt time delay model, we will not be able to measure the Hubble constant even with 100 GW detections. This is a consequence of the fact that the determination of $\hat{\alpha}_{0}$ is degenerate with the prompt time delay distribution, which is a crucial ingredient to get the implicit redshift information required for $H_{0}$. In other words, $H_{0}$ and the GW propagation speed are correlated with each other. The two parameters are degenerate since different combinations of their values can result in the same observed time delay. We display the correlations among the population parameters in Fig. 11, where we show the joint posterior distribution for 100 events detected by ET.